A 1.0-mW laser has a beam radius of 0.7mm . What is the intensity of this beam?
Answer #1

Intensity is power per unit area

area = A = area of the circle of radius r = 0.7 mm

A = \pi r^2 =\left ( \frac{22}{7} \right )*(0.7*10^{-3})^2 = 1.54*10^{-6} m^2

Intensity = \frac{Power}{Area}

Intensity = \frac{1*10^{^{-3}}}{1.54*10^{-6}} = 650 W/m^2